OpenStudy (anonymous):

If 6-foot-tall Madeleine casts an 108-inch shadow, how tall is her friend if, at the same time, his shadow is 1.5 feet shorter than hers?

OpenStudy (***[isuru]***):

the shadow of Madeleine in feet is = 108/12 = 9 the shadow of her friend = 9 - 1.5 = 7.5 now use the ratio... if the height of her friend is x then \[\frac{ 6 }{ x } = \frac{ 9 }{ 7.5 }\] where u will get the height of her friend as 5 feet hope this will help ya!!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (***[isuru]***):

u r welcome!